# Ultra-cold atoms as quantum simulators for relativistic phenomena

### v1

@@ Abstract
The goal of this article is to review developments regarding the use of ultra-cold atoms as quantum simulators. Special emphasis is placed on relativistic quantum phenomena, which are presumably most interesting for the audience of this journal. After a brief introduction into the main idea of quantum simulators and the basic physics of ultra-cold atoms, relativistic quantum phenomena of linear fields are discussed, including Hawking radiation, the Unruh effect, cosmological particle creation, the Gibbons-Hawking and Ginzburg effects, super-radiance, Sauter-Schwinger and Breit-Wheeler pair creation, as well as the dynamical Casimir effect. After that, the focus is shifted to phenomena of non-linear fields, such as the sine-Gordon model, the Kibble-Zurek mechanism, false-vacuum decay, and quantum back-reaction. In order to place everything into proper context, the basic underlying mechanisms of these phenomena are briefly recapitulated before their simulators are discussed. Even though effort is made to provide a review as fair as possible, there can be co claim of completeness and the selection as well as the relative weights of the topics may well reflect some personal bias.
@@ 1.2.2   Measurement schemes
Of course, realizing the Hamiltonian which allows us to simulate the phenomenon under consideration is not the end of the story, we also need a read-out scheme in order to measure the observables we are interested in. For ultra-cold atoms, the majority of the schemes relies on measuring the positions of the atoms [^3] However, alternative ideas are conceivable as well, see, e.g., [54, 55]. or the position-dependent number or density of atoms in some way, e.g., by taking pictures or by sending the atoms into an atom detector, see, e.g., [56]. Obviously, the former method can be less destructive than the latter and thus could be repeated several times during the course of evolution. However, one should keep in mind that such a non-destructive measurement is never “interaction-free”. According to the laws of quantum theory, such a measurement changes the quantum state of the system (unless it happens to be in an atom number eigen-state) and thus perturbs it. This perturbation can be small (e.g., in the case of weak measurements) but it cannot be avoided completely.
@@ 1.2.3   Hubbard Hamiltonians
Even though Hubbard Hamiltonians belong more to condensed-matter or solid-state physics than to relativistic particle physics, let us briefly discuss them – possible relations will hopefully become more evident below. Focusing on bosons for a moment, we may expand the field operators $\hat{\psi}_{a}(\mbox{\boldmath$r$})$ into a complete set of orthonormal basis functions $f_{a}^{I}(\mbox{\boldmath$r$})$ via
@@
which is even more famous than the bosonic version (1.7). Naturally, this system has also been realized experimentally, see, e.g., [57, 67, 68, 69] as well as [70, 71]. Due to the additional spin degree $s\in\{\uparrow,\downarrow\}$, its dynamics is more involved and less well understood. Nevertheless, for large $U$, the ground state does also become a Mott insulator in the case of half filling $\langle\hat{n}_{I}^{\uparrow}\rangle=\langle\hat{n}_{I}^{\downarrow}\rangle=1/2$. For small $U$, on the other hand, the state would be metallic (in more than one dimension), i.e., conducting. In the Mott insulator state, conduction requires the generation of quasi-particle excitations in the form of doublons (i.e., doubly occupied lattice sites) or holons (i.e., empty lattice sites). The proper description of their dynamics, including the interactions among each other (see, e.g., [72] and references therein) is an open question and might also be relevant for our understanding of high-temperature superconductivity.
@@ 1.2.5   Scalar field
To this end, we consider a chain (or array) of atoms with spin zero which are strongly localized in the $x$ and $y$ directions by suitable potentials (e.g., a deep optical lattice) such than they stay in the lowest quantum well states in those directions, see Fig. 1.5. Thus, we may focus on their motion in the remaining $z$ direction where we assume that we have one atom in each lattice site $I$. Denoting the position of the atom with mass $m$ in the lattice site $I$ by $q_{I}$, this motion in $z$ direction is influenced by the potential $V(q_{I})$ which is supposed to be much weaker than the confining potentials in the $x$ and $y$ directions. Furthermore, we assume that the lattice sites are close enough that we may have a strong coupling (e.g., by having Rydberg atoms) between the lattice sites $W(q_{I+1}-q_{I})$ where we neglect to the coupling between next to nearest neighbors, see Fig. 1.5. Together with the usual kinetic term, the Lagrangian for this system reads
@@
Figure 1.5: Sketch of the simulator for the scalar-field Lagrangian (1.22).
@@ 2   Linear fields
where $g_{s}$ denotes the contact ($s$-wave scattering) interaction strength. In a Bose-Einstein condensate, almost all atoms occupy the same quantum state as described by condensate wave function $\psi_{\rm c}(t,\mbox{\boldmath$r$})$ which can be time-dependent in general, see, e.g., [98]. Thus, we may apply the mean-field approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\psi_{\rm c}(t,\mbox{\boldmath$r$})$ where the many-particle field operator $\hat{\psi}(t,\mbox{\boldmath$r$})$ is effectively replaced by a c-number valued function $\psi_{\rm c}(t,\mbox{\boldmath$r$})$. Note that this approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\psi_{\rm c}(t,\mbox{\boldmath$r$})$ is a strongly simplified picture. For example, in a state with a definite number of atoms
@@
we would have $\langle\hat{\psi}\rangle=0$ which contradicts the mean-field approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\psi_{\rm c}(t,\mbox{\boldmath$r$})$. This contradiction can be repaired by a refined mean-field approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\hat{A}\hat{N}^{-1/2}\psi_{\rm c}(t,\mbox{\boldmath$r$})$ where $\hat{A}^{\dagger}$ and $\hat{A}$ are suitable creation and annihilation operators for the total number of atoms in the cloud $\hat{N}=\hat{A}^{\dagger}\hat{A}$, see, e.g., [99, 100, 101]. More refined expansion schemes such as $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\hat{A}\hat{N}^{-1/2}[\psi_{\rm c}(t,\mbox{\boldmath$r$})+\hat{\chi}(t,\mbox{\boldmath$r$})]$ can also include the quantum or thermal corrections $\hat{\chi}$ or fluctuations around this mean field or even higher-order corrections, see, e.g., [102].
@@
where the healing length $\xi=1/\sqrt{mg_{s}\rho_{\rm c}^{0}}$ can be used to separate the two regimes. For $\xi^{2}\mbox{\boldmath$k$}^{2}\ll 1$ we have the usual linear dispersion relation for sound $(\omega-\mbox{\boldmath$v$}_{\rm c}^{0}\cdot\mbox{\boldmath$k$})^{2}=c_{\rm s}^{2}\mbox{\boldmath$k$}^{2}$ while for $\xi^{2}\mbox{\boldmath$k$}^{2}\gg 1$, the dispersion relation becomes quadratic and approaches the free-particle form $(\omega-\mbox{\boldmath$v$}_{\rm c}^{0}\cdot\mbox{\boldmath$k$})^{2}=[\hbar^{2}\mbox{\boldmath$k$}^{2}/(2m)]^{2}$. As a result, phonons with large $k$ can move faster than $c_{\rm s}$, which will be important for the discussions in the next section.
@@ 2.2   Hawking radiation
where $M$ is the mass of the black hole. As a visualization of the order of magnitude, the typical wavelength of photons emitted at such a temperature is comparable to the Schwarzschild radius $R_{\rm S}=2G_{\rm N}M/c^{2}$ which marks the location of the horizon (blue curve in Fig. 2.1 left). Even though this temperature is extremely small for typical astronomical black holes, it is considered to be very important for our fundamental understanding of nature and has many ramifications such as the concept of black-hole entropy etc.
As an intuitive explanation for Hawking radiation, often the picture based on particle–anti-particle pairs is used. In this picture, the vacuum is visualized as containing many “virtual” particle–anti-particle pairs which are constantly created and annihilated. If one of them (the in-falling partner particle, see also [111]) is then trapped behind the horizon, the other one is free to escape to $r\to\infty$, i.e., it becomes “real” and constitutes the Hawking radiation. This intuitive picture may be appropriate for visualizing the quantum vacuum fluctuations and their entanglement at finite spatial distances (which are important ingredients for Hawking radiation), but one should be careful and not take it too seriously. For example, taking this picture too seriously, one has to entertain the possibility that a photon and an ”anti-photon” annihilate in vacuum without creating any new particles. Furthermore, the concept of classical trajectories for both, the outgoing Hawking radiation and the in-falling partner particle, which are supposed to originate from the same space-time point, does not quite work.
@@
Now, if we were to switch off the coupling between the two oscillators slowly, i.e., adiabatically, this entangled ground state would evolve into the factorizing ground state of the two individual oscillators discussed above. However, if the coupling is switched off rapidly, i.e., non-adiabatically (say suddenly), the quantum state has not enough time to react and the final state retains some or even all of the initial entanglement, i.e., it deviates from the ground state. In the case of Hawking radiation, this non-adiabatic change arises from the tearing apart of the waves at the horizon, where one part of the wave (represented by one oscillator) falls into the black hole while the other part of the wave (represented by the other oscillator) escapes to infinity. This tearing apart of waves as a non-adiabatic process causes a deviation from the ground state and its rapidity is set by the surface gravity of the black hole, which in turn determined the temperature (2.13). Even though the combined quantum state of the two oscillators is a pure state, the reduced state of the oscillator which escaped to infinity is a thermal, i.e., mixed state. In this picture, the two oscillators initially represent the quantum vacuum fluctuations at two different positions (which are also correlated [112]) which then evolve on trajectories on either side of the horizon, see Fig. 2.1. Due to the tearing apart by the gravitational field of the black hole, one of them escapes to infinity and represents the outgoing Hawking radiation (in a thermal state) while to other one corresponds to the in-falling partner [111]. For real black hole, the correlation between the outgoing Hawking particles and their in-falling partners can obviously not be observed, but for the black-hole analogues discussed below, this is an important point for the experimental observations, see also [113, 114].
@@
As a somewhat related issue, it should also be mentioned here that Hawking’s derivation [108, 109] is not without any problems. An important point is the so-called trans-Planckian problem [118, 119] which can be illustrated in the following way: If we consider an outgoing particle of the Hawking radiation far away from the black hole and trace it back in time, we have to undo the gravitational red-shift. Thus, its wave packet is squeezed against the horizon (when going back in time) and the wavelengths become shorter and shorter. Since the gravitational red-shift is exponential in time $t$ (where the exponent is determined to the surface gravity of the black hole), these wavelengths become smaller than the Planck length $\ell_{\rm Planck}=\sqrt{\hbar G_{\rm N}/c^{3}}\approx 1.6\times 10^{-35}~{}\rm m$ after a comparably short time. At this point, one would probably not trust the approach used by Hawking based on the propagation of quantum fields on a given classical background space-time anymore.
@@ 2.2.1   Black hole analogues
Actually, the trans-Planckian problem sketched above was one of the main motivations for Unruh [104] to consider laboratory analogues for black holes – or, more precisely, for Hawking radiation. His main idea is already explained in Sec. 2.1 and can be based on the quantitative analogy between Eqs. (2.10) and (2.1). To see how the analogue of a black hole and thus Hawking radiation can arise in such a laboratory set-up, let us consider a de Laval nozzle, see Fig. 2.2. At the entrance of the nozzle, the flow is sub-sonic $v_{\rm c}<c_{\rm s}$ and thus sound waves (which represent the analogue of light rays) can propagate in all directions (though dragged by the flow). This region corresponds to the black-hole exterior $r>R_{\rm S}$. At the exit of the nozzle, the flow is super-sonic $v_{\rm c}>c_{\rm s}$ such that sound waves cannot propagate upstream anymore and are swept away. This region corresponds to the black-hole interior $r<R_{\rm S}$. At the narrowest point of the nozzle, the flow velocity equals the speed of sound $v_{\rm c}=c_{\rm s}$ and thus this point is the analogue of the horizon $r=R_{\rm S}$.
@@
For Bose-Einstein condensates [120], we know what happens at short wavelengths, as explained in Sec. 2.1, and thus we can address the trans-Planckian problem by means of this analogue. In this case, the healing length $\xi$ where the dispersion relation (2.12) changes from linear to quadratic can be used as an analogue for the Planck length. Then the trans-Planckian problem is resolved in the following way: The initial wave packet has large wave number $k\gg 1/\xi$ and is thus faster than $c_{\rm s}$ such that it can propagate upstream, i.e., it originates in the super-sonic region. While it propagates upstream, it is stretched (as the analogue of the gravitational red-shift) because the flow velocity on its right-hand side is smaller than on its left-hand side, see Fig. 2.3. Around the time this wave packet reaches the horizon, its wave number is reduced to $k=\,{\cal O}(1/\xi)$ and thus its group velocity drops down to the sound speed. After that, the part of the wave packet which was already outside the horizon $r>R_{\rm S}$ can escape while the other part which was inside the horizon $r<R_{\rm S}$ is swept away. Hence, we find basically the same tearing apart process as discussed above. Actually, detailed calculations (see, e.g., [121, 122, 123] and references therein) show that – if we start with a wave packet in its initial ground state in the locally co-moving fluid frame (which is analogous to the freely falling frame in gravity) – the state of the part of the wave packet which escapes the nozzle on the right-hand side (i.e., its entrance) is indeed a thermal state with the analogue Hawking temperature

### v2

@@ Abstract
The goal of this article is to review developments regarding the use of ultra-cold atoms as quantum simulators. Special emphasis is placed on relativistic quantum phenomena, which are presumably most interesting for the audience of this journal. After a brief introduction into the main idea of quantum simulators and the basic physics of ultra-cold atoms, relativistic quantum phenomena of linear fields are discussed, including Hawking radiation, the Unruh effect, cosmological particle creation, the Gibbons-Hawking and Ginzburg effects, super-radiance, Sauter-Schwinger and Breit-Wheeler pair creation, as well as the dynamical Casimir effect. After that, the focus is shifted to phenomena of non-linear fields, such as the sine-Gordon model, the Kibble-Zurek mechanism, false-vacuum decay, and quantum back-reaction. In order to place everything into proper context, the basic underlying mechanisms of these phenomena are briefly recapitulated before their simulators are discussed. Even though effort is made to provide a review as fair as possible, there can be co claim of completeness and the selection as well as the relative weights of the topics may well reflect the personal view and taste of the author.
@@ 1.2.2   Measurement schemes
Of course, realizing the Hamiltonian which allows us to simulate the phenomenon under consideration is not the end of the story, we also need a read-out scheme in order to measure the observables we are interested in. For ultra-cold atoms, the majority of the schemes relies on measuring the positions of the atoms [^3] However, alternative ideas are conceivable as well, see, e.g., [54, 55]. or the position-dependent number or density of atoms in some way, e.g., by taking pictures or by sending the atoms into an atom detector, see, e.g., [56]. Obviously, the former method can be less destructive than the latter and thus could be repeated several times during the course of evolution. However, one should keep in mind that such a non-destructive measurement is never “interaction-free”. According to the laws of quantum theory, such a measurement changes the quantum state of the system (unless it happens to be in an atom number eigen-state) and thus perturbs it. This perturbation or back-action can be small (e.g., in the case of weak measurements), but it cannot be avoided completely.
@@ 1.2.3   Hubbard Hamiltonians
Even though the main focus of this review lies on relativistic particle physics, whereas Hubbard Hamiltonians belong more to condensed-matter or solid-state physics, let us briefly discuss them – possible relations will hopefully become more evident below. Focusing on bosons for a moment, we may expand the field operators $\hat{\psi}_{a}(\mbox{\boldmath$r$})$ into a complete set of orthonormal basis functions $f_{a}^{I}(\mbox{\boldmath$r$})$ via
@@
which is even more famous than the bosonic version (1.7). Naturally, this system has also been realized experimentally, see, e.g., [57, 67, 68, 69] as well as [70, 71]. Due to the additional spin degree of freedom $s\in\{\uparrow,\downarrow\}$, its dynamics is more involved and less well understood. Nevertheless, for large $U$, the ground state does also become a Mott insulator in the case of half filling $\langle\hat{n}_{I}^{\uparrow}\rangle=\langle\hat{n}_{I}^{\downarrow}\rangle=1/2$. For small $U$, on the other hand, the state would be metallic (in more than one dimension), i.e., conducting. In the Mott insulator state, conduction requires the generation of quasi-particle excitations in the form of doublons (i.e., doubly occupied lattice sites) or holons (i.e., empty lattice sites). The proper description of their dynamics, including the interactions among each other (see, e.g., [72] and references therein) is an open question and might also be relevant for our understanding of high-temperature superconductivity.
@@ 1.2.5   Scalar field
To this end, we consider a chain (or array) of atoms with spin zero which are strongly localized in the $x$ and $y$ directions by suitable potentials (e.g., a deep optical lattice) such than they stay in the lowest quantum well states in those directions, see Fig. 1.5. Thus, we may focus on their motion in the remaining $z$ direction where we assume that we have one atom in each lattice site $I$. Denoting the position of the atom with mass $m$ in the lattice site $I$ by $q_{I}$, this motion in $z$ direction is influenced by the potential $V(q_{I})$ which is supposed to be much weaker than the confining potentials in the $x$ and $y$ directions. Furthermore, we assume that the lattice sites are close enough that we may have a strong coupling (e.g., by having Rydberg atoms) between the lattice sites $W(q_{I+1}-q_{I})$ where we neglect to the coupling between next to nearest neighbors, see Fig. 1.5. Together with the usual kinetic energy term, the Lagrangian for this system reads
@@
Figure 1.5: Sketch of the simulator for the scalar-field Lagrangian (1.22). On the left-hand side, the atoms at the positions $q_{I}$ are displayed as blue dots together with their interactions $W$ between nearest neighbors. On the right-hand side, the local on-site potentials $V(q_{I})$ are depicted, where the axes are rotated in order to align the $q_{I}$ direction with that on the left-hand side.
@@ 2   Linear fields
where $g_{s}$ denotes the contact ($s$-wave scattering) interaction strength. In a Bose-Einstein condensate, almost all atoms occupy the same quantum state as described by condensate wave function $\psi_{\rm c}(t,\mbox{\boldmath$r$})$ which can be time-dependent in general, see, e.g., [98]. Thus, we may apply the mean-field approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\psi_{\rm c}(t,\mbox{\boldmath$r$})$ where the many-particle field operator $\hat{\psi}(t,\mbox{\boldmath$r$})$ is effectively replaced by a c-number valued function $\psi_{\rm c}(t,\mbox{\boldmath$r$})$. Note that this approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\psi_{\rm c}(t,\mbox{\boldmath$r$})$ is a strongly simplified picture. For example, considering a state with a definite number of atoms
@@
and taking the quantum expectation value of the field operator $\hat{\psi}$, we would find that it vanishes $\langle\hat{\psi}\rangle=0$ which contradicts the mean-field approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\psi_{\rm c}(t,\mbox{\boldmath$r$})$. This contradiction can be repaired by a refined mean-field approximation $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\hat{A}\hat{N}^{-1/2}\psi_{\rm c}(t,\mbox{\boldmath$r$})$ where $\hat{A}^{\dagger}$ and $\hat{A}$ are suitable creation and annihilation operators for the total number of atoms in the cloud $\hat{N}=\hat{A}^{\dagger}\hat{A}$, see, e.g., [99, 100, 101]. More refined expansion schemes such as $\hat{\psi}(t,\mbox{\boldmath$r$})\approx\hat{A}\hat{N}^{-1/2}[\psi_{\rm c}(t,\mbox{\boldmath$r$})+\hat{\chi}(t,\mbox{\boldmath$r$})]$ can also include the quantum or thermal corrections $\hat{\chi}$ or fluctuations around this mean field or even higher-order corrections, see, e.g., [102].
@@
where the healing [^6] The healing length corresponds to the distance over which the condensate wave function (or order parameter) can heal, i.e., recover from a local perturbation, such as a vortex. It can be determined by balancing the kinetic energy density $\hbar^{2}|\mbox{\boldmath$\nabla$}\psi_{\rm c}|^{2}/(2m)$ and the internal interaction energy density $g_{s}|\psi_{\rm c}|^{4}/2$. length $\xi=1/\sqrt{mg_{s}\rho_{\rm c}^{0}}$ can be used to separate the two regimes. For $\xi^{2}\mbox{\boldmath$k$}^{2}\ll 1$ we have the usual linear dispersion relation for sound $(\omega-\mbox{\boldmath$v$}_{\rm c}^{0}\cdot\mbox{\boldmath$k$})^{2}=c_{\rm s}^{2}\mbox{\boldmath$k$}^{2}$ while for $\xi^{2}\mbox{\boldmath$k$}^{2}\gg 1$, the dispersion relation becomes quadratic and approaches the free-particle form $(\omega-\mbox{\boldmath$v$}_{\rm c}^{0}\cdot\mbox{\boldmath$k$})^{2}=[\hbar^{2}\mbox{\boldmath$k$}^{2}/(2m)]^{2}$. As a result, phonons with large $k$ can move faster than $c_{\rm s}$, which will be important for the discussions in the next section.
@@ 2.2   Hawking radiation
where $M$ is the mass of the black hole. As a visualization of the order of magnitude, the typical wavelength of photons emitted at such a temperature is comparable to the Schwarzschild radius $R_{\rm S}=2G_{\rm N}M/c^{2}$ which marks the location of the horizon (blue curve in Fig. 2.1 left). Even though this temperature is extremely small for typical astronomical black holes, it is considered to be very important for our fundamental understanding of nature and has many ramifications such as the concept of black-hole entropy etc. Hawking’s discovery [108, 109] sparked an enormous amount of research activity and can be considered one of the most influential findings of modern fundamental physics.
As an intuitive explanation for Hawking radiation, often the picture based on particle–anti-particle pairs is used. In this picture, the vacuum is visualized as containing many “virtual” particle–anti-particle pairs which are constantly created and annihilated. If one of them (the in-falling partner particle, see also [111]) is then trapped behind the horizon, the other one is free to escape to $r\to\infty$, i.e., it becomes “real” and constitutes the Hawking radiation. This intuitive picture may be appropriate for visualizing the quantum vacuum fluctuations and their entanglement at finite spatial distances (which are important ingredients for Hawking radiation), but one should be careful and not take it too seriously. For example, taking this picture too seriously, one has to entertain the possibility that a photon and an ”anti-photon” annihilate in vacuum without creating any new particles. Furthermore, the concept of classical trajectories for both, the outgoing Hawking radiation and the in-falling partner particle, which are supposed to originate from the same space-time point, does not quite work, see also Fig. 2.1.
@@
Now, if we were to switch off the coupling between the two oscillators slowly, i.e., adiabatically, this entangled ground state would evolve into the factorizing ground state of the two individual oscillators discussed above. However, if the coupling is switched off rapidly, i.e., non-adiabatically (say suddenly), the quantum state has not enough time to react and the final state retains some or even all of the initial entanglement, i.e., it deviates from the ground state. In the case of Hawking radiation, this non-adiabatic change arises from the tearing apart of the waves at the horizon, where one part of the wave (represented by one oscillator) falls into the black hole while the other part of the wave (represented by the other oscillator) escapes to infinity. This tearing apart of waves as a non-adiabatic process causes a deviation from the ground state and its rapidity is set by the surface gravity of the black hole, which in turn determines the temperature (2.13). Even though the combined quantum state of the two oscillators is a pure state, the reduced state of the oscillator which escaped to infinity is a thermal, i.e., mixed state. In this picture, the two oscillators initially represent the quantum vacuum fluctuations at two different positions (which are also correlated [112]) which then evolve on trajectories on either side of the horizon, see Fig. 2.1. Due to the tearing apart by the gravitational field of the black hole, one of them escapes to infinity and represents the outgoing Hawking radiation (in a thermal state) while to other one corresponds to the in-falling partner [111]. For real black hole, the correlation between the outgoing Hawking particles and their in-falling partners can obviously not be observed, but for the black-hole analogues discussed below, this is an important point for the experimental observations, see also [113, 114].
@@
As a somewhat related issue, it should also be mentioned here that Hawking’s derivation [108, 109] is not without any problems. An important point is the so-called trans-Planckian problem [118, 119] which can be illustrated in the following way: If we consider an outgoing particle of the Hawking radiation far away from the black hole and trace it back in time, we have to undo the gravitational red-shift. Thus, its wave packet is squeezed against the horizon (when going back in time) and the wavelengths become shorter and shorter. Since the gravitational red-shift is exponential in time $t$ (where the exponent is determined by the surface gravity of the black hole), these wavelengths become smaller than the Planck length $\ell_{\rm Planck}=\sqrt{\hbar G_{\rm N}/c^{3}}\approx 1.6\times 10^{-35}~{}\rm m$ after a comparably short time. At this point, one would probably not trust the approach used by Hawking based on the propagation of quantum fields on a given classical background space-time anymore.
@@ 2.2.1   Black hole analogues
Actually, the trans-Planckian problem sketched above was one of the main motivations for Unruh [104] to consider laboratory analogues for black holes – or, more precisely, for Hawking radiation. His main idea is already explained in Sec. 2.1 and can be based on the quantitative analogy between Eqs. (2.10) and (2.1). To see how the analogue of a black hole and thus Hawking radiation can arise in such a laboratory set-up, let us consider a de Laval nozzle, see Fig. 2.2. At the entrance of the nozzle, the flow is sub-sonic $v_{\rm c}<c_{\rm s}$ and thus sound waves (which represent the analogue of light rays) can propagate in all directions (though dragged by the flow). This region corresponds to the black-hole exterior $r>R_{\rm S}$. At the exit of the nozzle, the flow is super-sonic $v_{\rm c}>c_{\rm s}$ such that sound waves cannot propagate upstream anymore and are swept away. This region corresponds to the black-hole interior $r<R_{\rm S}$. At the narrowest point of the nozzle [^11] These properties follow from the equations of fluid dynamics and do not require specific fine tuning. Actually, the somewhat counter-intuitive fact that the fluid flow, once it exceeds the speed of sound, speeds up further when the nozzle becomes wider again is used in rocket engines, for example., the flow velocity equals the speed of sound $v_{\rm c}=c_{\rm s}$ and thus this point is the analogue of the horizon $r=R_{\rm S}$.
@@
For Bose-Einstein condensates [120], we know what happens at short wavelengths, as explained in Sec. 2.1, and thus we can address the trans-Planckian problem by means of this analogue. In this case, the healing length $\xi$ where the dispersion relation (2.12) changes from linear to quadratic can be used as an analogue for the Planck length [^12] For other scenarios, the dispersion relation and thus the analogue of the Planck length can be different. E.g., dipolar Bose-Einstein condensates can have a dispersion relation with an additional “roton” dip, where the group velocity vanishes, see also [121, 122].. Then the trans-Planckian problem is resolved in the following way: The initial wave packet has large wave number $k\gg 1/\xi$ and is thus faster than $c_{\rm s}$ such that it can propagate upstream, i.e., it originates in the super-sonic region. While it propagates upstream, it is stretched (as the analogue of the gravitational red-shift) because the flow velocity on its right-hand side is smaller than on its left-hand side, see Fig. 2.3. Around the time this wave packet reaches the horizon, its wave number is reduced to $k=\,{\cal O}(1/\xi)$ and thus its group velocity drops down to the sound speed. After that, the part of the wave packet which was already outside the horizon $r>R_{\rm S}$ can escape while the other part which was inside the horizon $r<R_{\rm S}$ is swept away. Hence, we find basically the same tearing apart process as discussed above. Actually, detailed calculations (see, e.g., [123, 124, 125] and references therein) show that – if we start with a wave packet in its initial ground state in the locally co-moving fluid frame (which is analogous to the freely falling frame in gravity) – the state of the part of the wave packet which escapes the nozzle on the right-hand side (i.e., its entrance) is indeed a thermal state with the analogue Hawking temperature